OpenStudy (anonymous):

An artist sells earrings from a booth at a fair. rent for the booth is $320. the artist makes $15.00 from each pair of earrings sold. must the artist sell to earn a profit of $700?

OpenStudy (anonymous):

Let x be the number of pairs of earrings Revenues - Expenses = Profit 15x - 320 = 700

OpenStudy (anonymous):

how many pairs of earrings before must

OpenStudy (anonymous):

Solve 15x - 320 = 700 for x

OpenStudy (anonymous):

335

OpenStudy (anonymous):

help dnt get it

OpenStudy (anonymous):

15x - 320 = 700 Step one....add 320 to both sides. Now what equation do you have?

OpenStudy (anonymous):

wait do i put 320 under the 15x

OpenStudy (anonymous):

ADD..not divide..ADD 320 to both sides of the equation

OpenStudy (anonymous):

15x - 320 = 700 + 320 +320 15x = 1020

OpenStudy (anonymous):

Now, how do I get rid of the 15?

OpenStudy (anonymous):

Divide both sides of the equation by 15 15x = 1020 ---- ------ 15 15 x = 68 If you don't understand solving linear equation, you are REALLY going to be completely lost in all of your future math courses. Extremely important stuff. You will have to solve linear equations in every math course (and Physics and Chemistry and Economics etc) from heare on out. Go back and review until you get it.

OpenStudy (anonymous):

ok thx